### v1

@@ 3.2 Per-channel Quantization
The per-channel quantization aims to quantize the $i$th column $\mathbf{w}_{i}$ of $\mathbf{W}$ to $\delta_{i}\mathbf{q}_{i}$, where $\mathbf{q}_{i}$ is the bit-code vector. For the weight matrix $\mathbf{W}$, the quantized weight $\mathbf{W}_{q}\in\mathbb{R}^{m\times n}$ in per-channel quantization format is
$\mathbf{W}_{q}=\mathbf{Q}\text{diag}(\delta_{1},\ldots,\delta_{n})=(\delta_{1}\mathbf{q}_{1},\ldots,\delta_{n}\mathbf{q}_{n}).$
Thanks to (3), each column of $\mathbf{W}$ can be quantized independently by considering the following optimization
$\delta_{i},\mathbf{q}_{i}=\arg\min_{\delta,\mathbf{q}}\|\delta\mathbf{X}\mathbf{q}-\mathbf{X}\mathbf{w}_{i}\|^{2},\,i=1,\ldots,n.$ (8)
The COMQ method for per-layer quantization in Alg. 1 can be similarly applied to solving the above problem. Recall that the per-layer quantization uses the bit-code in the set $\mathbb{S}=\{-2^{b-1},\ldots,2^{b-1}-1\}$. This is because the whole weight matrix is centered around zero. However, the centers for each column can be different from zero. Therefore, we use the affine set $\mathbb{S}_{A}=\{z,z+1,\ldots,z+(2^{b}-1)\}$ for some proper zero-point $z$. In the following, we present the COMQ method to quantize the $i$-th column $\mathbf{w}_{i}$ into $b$-bit by solving (8).
Initialization. The scale factor is initialized as $\delta_{i}^{0}=\lambda\frac{\max(\mathbf{w}_{i})-\min(\mathbf{w}_{i})}{2^{b}-1}$ for some $0<\lambda\leq 1$. The $\lambda$ is to ensure we do not quantize most values to zero, especially for ultra-low bit quantization. The $\mathbf{q}_{i}^{0}$ is initialized as $\frac{\mathbf{w}_{i}}{\delta_{i}^{0}}$. The zero-point $z_{i}^{0}$ for the bit-code set is set as $z_{i}^{0}=\lfloor\frac{\min(\mathbf{w}_{i})}{\delta_{i}^{0}}\rceil$. The corresponding bit-code set is $\mathbb{S}_{z_{i}^{0}}=\{z_{i}^{0},z_{i}^{0}+1,\ldots,z_{i}^{0}+(2^{b}-1)\}$.
@@
Update. Let $\delta_{i}^{k-1},\mathbf{q}_{i}^{k-1}$ be the scale and quantized column produced by $(k-1)$-th iteration and the order of updating coordinates in $\mathbf{q}_{i}$ be $j_{1},\ldots,j_{m}$. After updating coordinates $j_{1},\ldots,j_{c-1}$, we update the coordinate $j_{c}$ as follows. We denote
$p_{j_{c}}=\frac{\left<\delta_{i}^{k-1}\mathbf{x}_{j_{c}},\mathbf{s}^{k}_{i,j_{c}}\right>}{\|\delta_{i}^{k-1}\mathbf{x}_{j_{c}}\|^{2}},$ (9)
where $\mathbf{s}^{k}_{i,j_{c}}=\mathbf{X}\mathbf{w}_{i}-\sum_{t\in\mathcal{J}_{-c}}\delta_{i}^{k-1}(\mathbf{q}_{i}^{k})_{t}\mathbf{x}_{t}-\sum_{t\in\mathcal{J}_{c}}\delta^{k-1}_{i}(\mathbf{q}_{i}^{k-1})_{t}\mathbf{x}_{t}$ for $\mathcal{J}_{-c}=\{j_{1},\ldots,j_{c-1}\}$ and $\mathcal{J}_{c}=\{j_{c+1},\ldots,j_{m}\}$. The value $p_{j_{c}}$ is the optimizer of (8) with respect to $(\mathbf{q}_{i})_{j_{c}}$ over the real field. Then, we project $p_{j_{c}}$ to the bit-code set $\mathbb{S}_{z_{i}^{k-1}}$, where $z_{i}^{k-1}=\lfloor\frac{\min(\mathbf{w}_{i})}{\delta_{i}^{k-1}}\rceil$. It is computed as
@@
$(\mathbf{q}_{i}^{k})_{j_{c}}=\text{clip}\left(\lfloor p_{j_{c}}\rceil,z_{i}^{k-1},z_{i}^{k-1}+2^{b}-1\right).$ (10)
After updating all coordinates in $\mathbf{q}_{i}$, we update the scale factor as
$\delta_{i}^{k}=\frac{\left<\mathbf{X}\mathbf{q}_{i}^{k},\mathbf{X}\mathbf{w}_{i}\right>}{\|\mathbf{q}_{i}^{k}\|^{2}}.$ (11)
Greedy update order. A popular choice for update order is the cyclic order, which updates the coordinates $(\mathbf{q}_{i})_{1},\ldots,(\mathbf{q}_{i})_{m}$ sequentially. However, it may not be optimal for per-channel quantization. To further reduce the quantization error and improve the performance of the quantized model, we propose a greedy update rule to determine the update order of the coordinates. The quantization error for the $i$th column can be written as
$\|\delta_{i}\mathbf{X}\mathbf{q}_{i}-\mathbf{X}\mathbf{w}_{i}\|^{2}=\left\|\sum_{j=1}^{m}\delta_{i}(\mathbf{q}_{i})_{j}\mathbf{x}_{j}-(\mathbf{w}_{i})_{j}\mathbf{x}_{j}\right\|^{2}.$
The importance of quantization target $(\mathbf{w}_{i})_{j}\mathbf{x}_{j}$ can be heuristically measured by its magnitude. Hence, we choose to first update the coordinates with a larger magnitude. This order ensures that the most significant coordinates are first updated and we have a sufficient quantization error decrease for each step. In practice, we sort $\|(\mathbf{w}_{i})_{1}\mathbf{x}_{1}\|,\ldots,\|(\mathbf{w}_{i})_{m}\mathbf{x}_{m}\|$ in descending order to get the coordinate update order $j_{1},\ldots,j_{m}$. To demonstrate the superiority of our proposed greedy order over the cyclic order, we compared the empirical layer-wise quantization errors $\|\mathbf{X}\mathbf{W}_{q}-\mathbf{X}\mathbf{W}\|$ across different architectures in Fig. 3.
@@
To summarize, the workflow for per-channel quantization is depicted in Fig. 2, and Alg. 2 summarizes the COMQ for layer-wise using per-channel PTQ.
Algorithm 2 Greedy COMQ for per-channel quantization
0:  A linear layer with weight $\mathbf{W}\in\mathbb{R}^{m\times n}$, input $\mathbf{X}$, and number of iterations $K$.
1:  for $i=1,\ldots,n$ do
2:     Initialize $\delta^{0}_{i},\mathbf{q}_{i}^{0}$ as described in Section 2
3:     Sort $\|(\mathbf{w}_{i})_{1}\mathbf{x}_{1}\|,\ldots,\|(\mathbf{w}_{i})_{m}\mathbf{x}_{m}\|$ in descending order to get update order $j_{1},\ldots,j_{m}$.
4:     for $k=1,\ldots,K$  do
5:        for $j=j_{1},\ldots,j_{m}$ do
6:           Find updated coordinate $(\mathbf{q}_{i}^{k})_{j}$ as in (10)
@@
8:        Find the new scale factor $\delta^{k}_{i}$ as in (11)
9:     end for
10:  end for
11:  Let $\mathbf{W}_{q}=(\delta_{1}^{K}\mathbf{q}_{1}^{K},\ldots,\delta_{n}^{K}\mathbf{q}_{n}^{K})$
11:  Quantized weight $\mathbf{W}_{q}$.
Method
Bit
ViT-S
ViT-B
DeiT-S
DeiT-B
Swin-T
Swin-S
Baseline
32/32
81.39
84.53
79.85
81.99
81.38
83.21
FQ-ViT [27]
4/32
81.09
76.23
79.92
78.81
81.89
PTQ4ViT [53]
4/32
72.34
72.06
77.50
80.07
78.46
80.24
Ours
4/32
80.35 83.86 78.98 81.40 80.89 82.85
FQ-ViT [27]
3/32
34.31
51.06
65.64
65.38
71.88
PTQ4ViT [53]
3/32
38.77
19.85
70.22
75.42
70.74
73.46
Ours
3/32
77.08 81.73 77.47 80.47 79.31 81.95
Ours
2/32
52.44 65.69 67.19 77.14 74.05 78.02
@@
Method
Bit
ViT-S
ViT-B
DeiT-S
DeiT-B
Swin-S
BRECQ [24]
4/4
12.36
9.68
63.73
72.31
72.74
QDrop [47]
4/4
21.42
47.30
68.27
72.60
79.58
APQ-ViT [5]
4/4
47.95
41.41
43.55
67.48
77.15
RepQ-ViT [25]
4/4
65.05
68.48
69.03
75.61
79.45
Ours
4/4
71.47 78.27 72.17 78.72 81.19
Ours
2/4
30.11 45.33 53.20 71.90 75.37
@@ 4.2.1 Vision Transformer
Method Bit ResNet18 ResNet50 MobileNetV2
@@ 4.2.2 Convolutional Neural Networks
Per-layer Quantization. Per-layer quantization is more computationally efficient but per-layer quantization in existing methods typically leads to high performance degradation. In contrast, our COMQ for per-layer quantization in Alg. 1 yields promising results, particularly on ResNets. With the float models from PyTorch, we compare COMQ with state-of-the-art uniform per-layer PTQ methods, such as AdaQuant [31] and Bit-split [43]. The results are presented in Table 3. For $4$-bit, COMQ achieves a mere 1.0% accuracy drop in ResNet18 and a 0.94% accuracy drop in ResNet50, outperforming other methods for per-layer PTQ. More remarkably, for $3$-bit quantization, COMQ achieves a 4.13% and an 7.9% accuracy drop on ResNet18 and ResNet50, respectively, surpassing all competing methods.
Method Bit ResNet18 ResNet50
@@
Method Bit ResNet18 ResNet50
@@
Table 6: Accuracy vs Batchsize for 4W32A per-channel PTQ.

### v3

@@ 3.2 Per-channel Quantization
The per-channel quantization aims to quantize the $j$-th column $\boldsymbol{w}_{i}$ of $\boldsymbol{W}$ to $\delta_{j}\boldsymbol{q}_{j}$, where $\boldsymbol{q}_{j}$ is bit-code vector. For the weight matrix $\boldsymbol{W}$, the quantized weight $\boldsymbol{W}_{q}\in\mathbb{R}^{m\times n}$ in per-channel quantization format is
$\boldsymbol{W}_{q}=\boldsymbol{Q}\text{diag}(\boldsymbol{\delta})=\boldsymbol{Q}\text{diag}(\delta_{1},\ldots,\delta_{n})=(\delta_{1}\boldsymbol{q}_{1},\ldots,\delta_{n}\boldsymbol{q}_{n}).$
Same as for per-layer quantization, the $j$-th column of $\boldsymbol{W}$ can be quantized independently by solving the following optimization problem:
$\delta_{j},\boldsymbol{q}_{j}=\arg\min_{\delta,\boldsymbol{q}}\|\delta\boldsymbol{X}\boldsymbol{q}-\boldsymbol{X}\boldsymbol{w}_{j}\|^{2},\,j\in[n].$ (8)
Initialization. The scale factor is initialized as $\delta_{j}^{0}=\lambda\frac{\max(\boldsymbol{w}_{j})-\min(\boldsymbol{w}_{j})}{2^{b}-1}$ for some $0<\lambda\leq 1$. The $\lambda$ is to ensure we do not quantize most values to zero, especially for ultra-low bit quantization. The $\boldsymbol{q}_{j}^{0}$ is initialized as $\frac{\boldsymbol{w}_{j}}{\delta_{j}^{0}}$.
@@
$\boldsymbol{Q}$-Update. Let $\boldsymbol{\delta}^{k-1}\in\mathbb{R}^{n}$ and $\boldsymbol{Q}^{k-1}\in\mathbb{S}^{m\times n}$ be the scaling factors and bit-code matrix produced by $(k-1)$-th iteration. Or equivalently, suppose $\boldsymbol{W}_{q}^{k-1}=\boldsymbol{Q}^{k-1}\text{diag}(\boldsymbol{\delta}^{k-1})$ is the current quantized weight matrix. The $\boldsymbol{Q}$-Update for per-channel quantization is substantially similar to the per-layer setting. The updated bit-code $Q_{i,j}^{k}$ is given by
@@
$Q^{k}_{i,j}=\text{clip}\left(\left\lfloor\frac{\left<\boldsymbol{x}_{i},\boldsymbol{s}^{k}_{i,j}\right>}{\delta_{j}^{k-1}\|\boldsymbol{x}_{i}\|^{2}}\right\rceil,z_{j},z_{j}+2^{b}-1\right),$
where $\boldsymbol{s}_{i,j}^{k}=\boldsymbol{X}\boldsymbol{w}_{j}-\delta_{j}^{k-1}(\sum_{t=1}^{i-1}Q_{t,j}^{k}\boldsymbol{x}_{t}+\sum_{t=i+1}^{m}Q_{t,j}^{k-1}\boldsymbol{x}_{t})$, and $z_{j}$ is the zero-point for quantizing the $j$-th column. For the row-wise update of $\boldsymbol{Q}^{k}$, we simply follow (3.1) with minor adaptions to the column-wise scaling. Using the same notations, we denote by $\boldsymbol{w}_{i,:}\in\mathbb{R}^{n}$ and $\boldsymbol{q}_{i,:}\in\mathbb{R}^{n}$ the $i$-th row of $\boldsymbol{W}$ and $\boldsymbol{Q}$, respectively, and denote by $\boldsymbol{x}_{:,i}\in\mathbb{R}^{m}$ the $i$-th column of $\boldsymbol{X}$. With $\boldsymbol{U}_{0}^{k}=\boldsymbol{X}(\boldsymbol{W}-\boldsymbol{W}_{q}^{k-1})$, we iterate for $i=1,\dots,m$:
$\boldsymbol{U}_{i}^{k}=\boldsymbol{U}_{i-1}^{k}-\boldsymbol{x}_{:,i}\otimes(\boldsymbol{w}_{i,:}-\boldsymbol{\delta}^{k-1}\odot\boldsymbol{q}_{i,:}^{k-1})$
$\boldsymbol{q}^{k}_{i,:}=\text{clip}\left(\left\lfloor\frac{(\boldsymbol{U}^{k}_{i}+\boldsymbol{x}_{:,i}\otimes\boldsymbol{w}_{i,:})^{\top}\boldsymbol{x}_{:,i}}{\boldsymbol{x}_{:,i}^{\top}\boldsymbol{x}_{:,i}}\oslash\boldsymbol{\delta}^{k-1}\right\rceil,\boldsymbol{z},\boldsymbol{z}+2^{b}-1\right)$ (9)
$\boldsymbol{U}_{i}^{k}=\boldsymbol{U}_{i}^{k}+\boldsymbol{x}_{:,i}\otimes(\boldsymbol{w}_{i,:}-\boldsymbol{\delta}^{k-1}\odot\boldsymbol{q}_{i,:}^{k}),$
where $\boldsymbol{U}_{i}^{k}$ maintains the quantization residuals and $\boldsymbol{z}$ is the vector of zero-points across all columns.
$\boldsymbol{\delta}$-Update. Having obtained $\boldsymbol{Q}^{k}$, we update the scaling factors as
$\delta^{k}_{j}=\frac{\left<\boldsymbol{X}\boldsymbol{q}_{j}^{k},\boldsymbol{X}\boldsymbol{w}_{j}\right>}{\|\boldsymbol{X}\boldsymbol{q}_{j}^{k}\|^{2}}.$ (10)
@@
To summarize, the workflow for per-channel quantization is depicted in Fig. 2, and Alg. 2 describes COMQ for layer-wise PTQ with per-channel scaling.
Algorithm 2 COMQ for the per-channel quantization of one linear layer
0:  Pre-trained weights $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, feature matirx $\boldsymbol{X}$, and iteration number $K$.
1:  Initialize $\boldsymbol{\delta}^{0},\boldsymbol{Q}^{0}$ as described in Section 2
2:  for $k=1,\ldots,K$  do
3:     for $i=1,\ldots,m$ do
4:        Update the coordinates $\{Q_{i,j}^{k}\}_{j\in[n]}$ or $\boldsymbol{q}^{k}_{i,:}$ as in (3.2)
5:     end for
6:     Update the scaling factors $\boldsymbol{\delta}^{k}$ as in (10)
@@
8:  Compute $\boldsymbol{W}_{q}=(\delta_{1}^{K}\boldsymbol{q}_{1}^{K},\ldots,\delta_{n}^{K}\boldsymbol{q}_{n}^{K})$
8:  Quantized weight $\boldsymbol{W}_{q}$.
3.3 Greedy COMQ
Coordinate descent update the coordinates of a vector in the cyclic order (or index order) by default, which may not be optimal. To further reduce the quantization error and improve the performance of the quantized model, we propose a greedy update rule to determine the update order of the coordinates. The quantization error for a column with weight vector $\boldsymbol{w}$ can be written as $\|\delta\boldsymbol{X}\boldsymbol{q}-\boldsymbol{X}\boldsymbol{w}\|^{2}=\left\|\sum_{i=1}^{m}\delta q_{i}\boldsymbol{x}_{i}-w_{i}\boldsymbol{x}_{i}\right\|^{2}.$ The importance of quantization target $w_{i}\boldsymbol{x}_{i}$ can be heuristically measured by its magnitude. Hence, we choose to first update the coordinates with a larger magnitude. This order ensures that the most significant coordinates are first updated, and we have a sufficient quantization error decrease for each step. In practice, we sort $\|w_{1}\boldsymbol{x}_{1}\|,\ldots,\|w_{m}\boldsymbol{x}_{m}\|$ in descending order to get the coordinate update order $i_{1},\ldots,i_{m}$, permuting the index set $[m]$. Therefore, the greedy COMQ for quantizing the weight matrix $\boldsymbol{W}$ requires column-wise sorting of $(\boldsymbol{v}\odot|\boldsymbol{w}_{1}|,\dots,\boldsymbol{v}\odot|\boldsymbol{w}_{n}|)$, where $\boldsymbol{v}=(\|\boldsymbol{x}_{1}\|,\dots,\|\boldsymbol{x}_{m}\|)^{\top}$ is the column-wise Euclidean norm of $\boldsymbol{X}$ in descending order. Then, we permute the columns of $\boldsymbol{W}$ and rows of $\boldsymbol{X}$ individually according to the sorted indices, followed by the quantization process. After that, the inverse permutations are performed on $\boldsymbol{W}$ and $\boldsymbol{X}$ to restore the original index order.
To demonstrate the superiority of our proposed greedy order over the cyclic order, we compared the empirical layer-wise quantization errors $\|\boldsymbol{X}\boldsymbol{W}_{q}-\boldsymbol{X}\boldsymbol{W}\|$ across different architectures in Fig. 3. Additionally, we also compared their performance on ImageNet in Tab. 8 in the Appendix.
@@ 3.3 Greedy COMQ
Method WBit ViT-S ViT-B DeiT-S DeiT-B Swin-T Swin-S
Baseline 32 81.39 84.53 79.85 81.99 81.38 83.21
FQ-ViT [29] 4 - 81.09 76.23 79.92 78.81 81.89
PTQ4ViT [56] 4 72.34 72.06 77.50 80.07 78.46 80.24
Ours 4 80.35 83.86 78.98 81.40 80.89 82.85
FQ-ViT [29] 3 - 34.31 51.06 65.64 65.38 71.88
PTQ4ViT [56] 3 38.77 19.85 70.22 75.42 70.74 73.46
Ours 3 77.08 81.73 77.47 80.47 79.31 81.95
Ours 2 52.44 65.69 67.19 77.14 74.05 78.02
@@
Method Bit (W/A) ViT-S ViT-B DeiT-S DeiT-B Swin-S
BRECQ [26] 4/4 12.36 9.68 63.73 72.31 72.74
QDrop [48] 4/4 21.42 47.30 68.27 72.60 79.58
APQ-ViT [7] 4/4 47.95 41.41 43.55 67.48 77.15
RepQ-ViT [27] 4/4 65.05 68.48 69.03 75.61 79.45
Ours 4/4 71.47 78.27 72.17 78.72 81.19
Ours 2/4 30.11 45.33 53.20 71.90 75.37
@@ 4.2.1 Vision Transformer
Method WBit ResNet18 ResNet50 MobileNetV2
@@ 4.2.2 Convolutional Neural Networks
Per-layer Quantization. Per-layer quantization is more computationally efficient but per-layer quantization in existing methods typically leads to high performance degradation. In contrast, our greedy COMQ for per-layer quantization in Alg. 1 yields promising results, particularly on ResNets. With the float models from PyTorch, we compare COMQ with state-of-the-art uniform per-layer PTQ methods, such as AdaQuant [33] and Bit-split [45]. The results are presented in Table 3. For $4$-bit, COMQ achieves a mere 1.0% accuracy drop in ResNet18 and a 0.94% accuracy drop in ResNet50, outperforming other methods for per-layer PTQ. More remarkably, for $3$-bit quantization, COMQ achieves a 4.13% and an 7.9% accuracy drop on ResNet18 and ResNet50, respectively, surpassing all competing methods. The cyclic COMQ for per-layer quantization is highlighted by a symbol.
Method Bit (W/A) ResNet18 ResNet50
@@
Method Bit (W/A) ResNet18 ResNet50
@@
Table 7: Accuracy vs Iteration Number $K$ for 4W32A per-layer PTQ.
@@ 5 Concluding Remarks
Acknowledgement
PY and AZ were partially supported by NSF grants DMS-2208126 and IIS-2110546. ZY was supported by NSF grant DMS-2110836 and a start-up grant from SUNY Albany. JX was supported by NSF grants DMS-2151225 and DMS-2219904, and a Qualcomm faculty gift award. XL was supported by NSF grant CCSS-2348046.
Appendix: More Ablation Study
We compare greedy update order and cyclic update order for the per-channel quantization in Algorithm 2. We evaluate the two update orders on five widely used models: ResNet18, ResNet50, ViT-S, DeiT-S, and Swin-T. The pre-trained ResNet family models are sourced from the PyTorch platform, and the pre-trained vision transformer models are trained on ImageNet using the open-source Timm library. The results are shown in Table 8. We can find that the greedy update order outperforms the cyclic update order in all test cases across all models and precisions, demonstrating that the proposed greedy update order greatly improves the performance of quantized models. Furthermore, the performance improvement is more significant for larger models at lower bit-widths.
Table 8: ImageNet results for cyclic and greedy COMQ with weight quantization.
Method Bits RN18 RN50 ViT-S DeiT-S Swin-T
FP 32 71.00 76.63 81.39 81.99 81.38
Cyclic 4 70.71 76.29 80.16 78.94 80.85
Greedy 70.83 76.38 80.35 78.98 80.89
Cyclic 3 69.53 75.58 76.58 77.20 78.81
Greedy 69.63 75.73 77.08 77.47 79.31
Cyclic 2 64.24 69.21 49.27 65.46 73.05
Greedy 64.52 70.32 52.44 67.19 74.05
The following Table 9 shows the runtime comparison on ResNet50 for the vector-wise update.
Table 9: Runtime of 4W32A COMQ (greedy), which is faster than the gradient-based methods.
Model AdaRound BRECQ OBQ COMQ
ResNet50 55 min 53 min 65 min 12 min
Table 10 shows the performance of our algorithm on Swin ViTs for different values of $\lambda$ in 2 bits. It is clear that the results of $lambda=0.71$ (near-optimal $\lambda$) are much better than that of $\lambda=1$.
Table 10: ImageNet accuracy for different $\lambda$ initialization. $\lambda=0.71$ is empirically (near-)optimal for Swin ViTs.
$\lambda$ Bits Swin-T Swin-S
1 2 65.05 70.06
0.71 74.05 78.02
FP 32 81.38 83.21